OpenStudy (swissgirl):

Is this a linear order? V where mVn iff m is odd and n is even, or m and n are even and \( m \leq n\) m and n are odd and \( m \leq n\)

OpenStudy (kinggeorge):

Integers or naturals?

OpenStudy (swissgirl):

Naturals

OpenStudy (swissgirl):

Like I am gonnnnaaa cry i cant solve a thing

OpenStudy (swissgirl):

all \( \leq \) in N are lineared ordered

OpenStudy (kinggeorge):

First up, we need to show antisymmetry. Suppose \(mVn\) and \(nVm\). Note that this means they are both either odd or both even. WLOG, suppose they are both even. So \(m\le n\) and \(n\le m\). Clearly, \(m=n\). So this is antisymmetric.

OpenStudy (swissgirl):

okkkkk

OpenStudy (kinggeorge):

Now, onto transitivity. Suppose \(rVs\) and \(sVt\). We have a few more cases to consider here. If they are all even/all odd we can see it's transitive because \(\le\) is transitive in the natural numbers. Now let's start working through the cases. Suppose \(r\) is even. Then \(s\) must be even, and so must \(t\) (do you see why?), so we've reduced the problem to what I wrote above. Suppose \(r\) is odd. Suppose \(s\) is then even, forcing \(t\) to be even. Since \(r\) is odd, and \(t\) is even, this is also transitive. Now suppose \(r\) is odd, \(s\) is odd, and \(t\) is either odd or even. If \(t\) is even, it's transitive by the first condition. If \(t\) is odd, it's transitive by the third condition and transitivity of \(\le\). Therefore, this is transitive.

OpenStudy (swissgirl):

AWESOME :)))))))))

OpenStudy (kinggeorge):

Now, we finally need to show totality of the ordering. Let \(a,b\in\mathbb{N}\) be arbitrary. We have 4 cases here, although it can be simplified WLOG to 2 cases. Case 1: \(a,b\) are both odd/even. Since \(\le\) is a linear order on \(\mathbb{N}\), we know that either \(a\le b\) or \(b\le a\). Case 2: One of \(a,b\) is odd, and the other is even. WLOG, suppose \(a\) is even, and \(b\) is odd. Then \(bVa\). Hence, this relation has totality, and is a linear order.

OpenStudy (swissgirl):

hahahahha still trying to follow the def of totality lol

OpenStudy (swissgirl):

How do u see totality with the second case?

OpenStudy (kinggeorge):

Totality is merely: Given \(a,b\), then \(aVb\) or \(bVa\) (or both). In the second case, there's really two cases there. Let me break it up a bit further.

OpenStudy (kinggeorge):

Case 2a: \(a\) is even, \(b\) is odd. Then \(bVa\), so we're done. Case 2b: \(b\) is even, \(a\) is odd. Then \(aVb\), so we're done. No matter what the case is, you'll always find that either \(aVb\) or \(bVa\), and so it has totality.

OpenStudy (swissgirl):

sorry just gonna ask one more question

OpenStudy (swissgirl):

but the first condition is symmetric isnt it?

OpenStudy (kinggeorge):

antisymmetric.

OpenStudy (swissgirl):

lol if a m is odd and n is even or m is odd or n is even

OpenStudy (swissgirl):

did u get what i was saying?

OpenStudy (kinggeorge):

No, sorry. Could you try and explain again? :/

OpenStudy (swissgirl):

hahahah okkkk

OpenStudy (swissgirl):

Like the first condition of this relation is m is odd and n is even

OpenStudy (swissgirl):

Thats mVn right

OpenStudy (swissgirl):

now like what wld nVm be?

OpenStudy (swissgirl):

I am just referring to the first condition

OpenStudy (kinggeorge):

If m is odd, and n is even, then \(n\cancel{V}m\). \((n,m)\) is not actually in V.

OpenStudy (swissgirl):

okkk gotchhhaaaaaa

OpenStudy (swissgirl):

Thanks u r awesomeeee

OpenStudy (swissgirl):

U r seriously saving me here

OpenStudy (kinggeorge):

No problem :P

OpenStudy (swissgirl):

TTTHHANNNKKSSSSSS :D

OpenStudy (kinggeorge):

You're welcome. :)